CONSTITUTION OF LINUX USERS GROUP (SINGAPORE)

NAME

1. This Society shall be known as the “Linux Users’ Group (Singapore)”, hereinafter referred to as the “Society”.

PLACE OF BUSINESS

2. Its place of business shall be at “360 Orchard Road #09-01A, International Building, Singapore 0923” or such other address as may subsequently be decided upon by the Executive Committee (hereinafter referred to as Committee), and approved by the Registrar of Societies. The Society shall carry out its activities only in places and premises which have the prior written approval from the relevant authorities, where necessary.

OBJECTS

3. The objects of the Society shall be to promote the growth and development of knowledge with respect to the development, evolution, use and popularity of the Linux Operating System and other associated software and hardware.

MEMBERSHIP

4. Membership is open to all persons who are interested in using the Linux Operating System.

5. Only members who are above 18 years of age shall have the right to vote and to hold office in the Society.

6. Category

The membership of the Society shall consist of Ordinary members, and Student members of the Society. Only Ordinary members shall have the right to vote and to hold office in the Society.

7. Eligibility

(a) Ordinary Members:

i. Ordinary membership is open to all persons.

(b) Student Members:

i. Persons who are pursuing full-time courses at any institution in Singapore shall be eligible for Student membership.

PROCEDURE OF ADMISSION

8. A person wishing to join the Society should submit his particulars to the Secretary on a prescribed form. The Committee shall have the power to approve the application or reject it without assigning any reason therefore. A copy of the constitution shall be furnished to every approved member.

ENTRANCE FEES, SUBSCRIPTIONS AND OTHER DUES

9. An entrance fee as follows:

Ordinary Member $30
Student Member $20

is payable within two weeks of election to membership, in default of which membership may be cancelled by order of the Committee.

10. Annual subscriptions are payable as follows:

Ordinary Member $60 per annum
Student Member $40 per annum

11. For retiree members, the entrance fee is waived and the annual subscription is reduced by half.

12. Annual subscriptions are payable in advance on the first day of January. If a member falls into arrears with his subscription or other dues, he shall be informed immediately by the Treasurer. If he fails to settle his arrears within 4 weeks of their becoming due, the President may order that his name be posted on the Society’s notice board and that he be denied the priveledges of membership until he settles his account. If he falls into arrears for more that 3 months, he will automatically cease to be a member and the Committee may take legal action against him provided that they are satisfied that he has received due notice of his debts.

13. The amount of fees may only be varied by the General Meeting of members. Any additional fund required for special purposes may only be raised from members with the consent of the General Meeting of the members.

SUPREME AUTHORITY AND GENERAL MEETINGS

14. The supreme authority of the Society is vested in a General Meeting of the members presided over by the President.

15. An Annual General Meeting shall be held in June every year.

16. At other times, an Extraordinary General Meeting must be called by the President on the request in writing of not less than one-quarter of the total voting membership or 30 voting members, whichever is the lesser, and may be called anytime by order of the Committee. The notice in writing shall be given to the Secretary setting forth the business that is to be transacted. The Extraordinary General Meeting shall be convened within two months from receiving this request to convene the Extraordinary General Meeting.

17. If the Committee does not within two months after the date of the receipt of the written request proceed to convene an Extraordinary General Meeting, the members who requested the Extraordinary General Meeting shall convene the Extraordinary General Meeting by giving 10 days’ notice to voting members setting forth the business to be transacted and simultaneously posting the agenda on the Society’s notice board.

18. At least two weeks’ notice shall be given of an Annual General Meeting and at least 10 day’s notice of an Extraordinary General Meeting. Notice of meeting stating the date, time and place of meeting shall be sent by the Secretary to all voting members. The particulars of the agenda shall be posted on the Society’s notice board four days in advance of the meeting.

19. The following points will be considered at the Annual General Meeting:

(a) The previous financial year’s accounts and annual report of the Committee.

(b) Where applicable, the election of office-bearers and Honorary Auditors for the following term.

20. Any member who wishes to place an item on the agenda of a General Meeting may do so provided he gives notice to the Secretary one week before the meeting is due to be held.

21. At least one-quarter of the total voting membership or 30 voting members, whichever is the lesser, present at a General Meeting shall form a quorum.

22. In the event of there being no quorum at the commencement of a General Meeting, the meeting shall be adjourned for half an hour and should the number then present be insufficient to form a quorum, those present shall be considered a quorum, but they shall have no power to amend any of the existing Rules.

MANAGEMENT AND COMMITTEE

23. The administration of the Society shall be entrusted to a Committee consisting of the following to be elected at each Annual General Meeting:

A President A Vice-President A Secretary An Assistant Secretary A Treasurer An Assistant Treasurer

24. Names for the above offices shall be proposed and seconded at the Annual General Meeting and election will follow on a simple majority vote of the members. All office-bearers, except the Treasurer and Assistant Treasurer, may be re-elected to the same or related post for a consecutive term of office. The term of office of the Committee is one year.

25. Election will be either by show of hands or, subject to the agreement of the majority of the voting members present, by secret ballots. In the event of a tie, the Chairman of the meeting shall have a casting vote.

26. A Committee Meeting shall be held at least once a month after giving seven days’ notice to Committee Members. The President may call a Committee Meeting at any time by giving five days’ notice. At least half of the Committee members must be present for its proceedings to be valid.

27. Any member of the Committee absenting himself from three meetings consecutively without satisfactory explanations shall be deemed to have withdrawn from the Committee and a successor may be co-opted by the Committee to serve until the next Annual General Meeting. Any change in the Committee shall be notified to the Registrar of Societies within two weeks of the change.

28. The duty of the Committee is to organise and supervise the daily activities of the Society. The Committee may not act contrary to the expressed wishes of the General Meeting without prior reference to it and always remains subordinate to the General Meeting.

29. The Committee has power to authorize the expenditure of a sum not exceeding $6000 per month from the Society’s funds for the Society’s purposes.

DUTIES OF OFFICE-BEARERS

30. The President shall act as Chairman at all General and Committee Meetings. He shall also represent the Society in its dealings with outside persons.

31. The Vice-President shall assist the President and deputise for him in his absence.

32. The Secretary shall keep all records, except financial, of the Society and shall be responsible for their correctness. He will keep minutes of all General and Committee Meetings. He shall maintain an up-to-date Registir of Members at all times.

33. The Assistant Secretary shall assist the Secretary and deputise for him in his absence.

34. The Treasurer shall keep all funds and collect and disburse all moneys on behalf of the Society and shall keep an account of all monetary transactions and shall be responsible for their correctness. He is authorised to expend up to $1000 per month for petty expenses on behalf of the Society. He will not keep more than $1000 in the form of cash and money in excess of this will be deposited in a bank to be named by the Committee. Cheques, etc for withdrawals from the bank will be signed by the Treasurer and either the President or the Vice-President or the Secretary.

35. The Assistant Treasurer shall assist the Treasurer and deputise for him in his absence.

AUDIT AND FINANCIAL YEAR

36. Two voting members, not being members of the Committee, shall be elected as Honorary Auditors at each Annual General Meeting and will hold office for a term of one year only and shall not be re-elected for a consecutive term. They:

(a) Will be required to audit each year’s accounts and present a report upon them to the Annual General meeting.

(b) May be required by the President to audit the Society’s accounts for any period within their tenure of office at any date and make a report to the Committee.

37. The financial year shall be from 1st April to 31st March.

PROHIBITIONS

38. (a) Gambling of any kind, whether for stakes or not, is forbidden on the Society’s premises. The introduction of materials for gambling or drug taking and of bad characters into the premises is prohibited.

(b) The funds of the Society shall not be used to pay the fines of members who have been convicted in Court.

(c) The Society shall not engage in any trade union activity as defined in any written law relating to trade unions for the time being in force in Singapore.

(d) The Society shall not attempt to restrict or interfere with trade or make directly or indirectly any recommendation to, any arrangement with its members which has the purpose or is likely to have the effect of fixing or controlling the price or any discount, allowance or rebate relating to any goods or services which adversely affect consumer interests.

(e) The Society shall not indulge in any political activity or allow its funds and/or premises to be used for political purposes.

(f) The Society shall not hold any lottery, whether confined to its members or not, in the name of the Society or its office-bearers, Committee or members unless with the prior approval of the relevant authorities.

(g) The Society shall not raise funds from the public for whatever purposes without the prior approval in writing of the Director, Criminal Investigation Department and other relevant authorities.

AMENDMENTS TO RULES

39. No alteration or addition/deletion to these rules shall be made except at a General Meeting and with the consent of a simple majority of the voting members present at the General Meeting and they shall not come into force without the prior sanction of the Registrar of Societies.

INTERPRETATION

40. In the event of any question or matter pertaining to day-to-day administration which is not expressly provided for in the rules, the Committee shall have power to use their own discretion. The decision of the Committee shall be final unless it is reversed at a General Meeting of members.

DISPUTES

41. In the event of any dispute arising amongst members, they shall attempt to resolve the matter at an Extraordinary General Meeting in accordance with the rules in the constitution. Should the members fail to resolve the matter, they may bring the matter to a court of law for settlement.

DISSOLUTION

42. (a) The Society shall not be dissolved, except with the consent of not less than three-fifths of the total voting membership of the Society for the time being resident in Singapore expressed, either in person or by proxy, at a General Meeting convened for the purpose.

(b) In the event of the Society being dissolved as provided above, all debts and liabilities legally incurred on behalf of the Society shall be fully discharged, and the remaining funds will be disposed of in such a manner as the General Meeting of members may determine or donated to an approved charity or charities in Singapore.

(c) A Certificate of Dissolution shall be given within seven days of the dissolution to the Registrar of Societies.